# LLaMA-Berry: Pairwise Optimization for O1-like Olympiad-Level Mathematical Reasoning

### v1

@@ Abstract
This paper presents an advanced mathematical problem-solving framework, LLaMA-Berry, for enhancing the mathematical reasoning ability of Large Language Models (LLMs). The framework combines Monte Carlo Tree Search (MCTS) with iterative Self-Refine to optimize the reasoning path and utilizes a pairwise reward model to evaluate different paths globally. By leveraging the self-critic and rewriting capabilities of LLMs, Self-Refine applied to MCTS (SR-MCTS) overcomes the inefficiencies and limitations of conventional step-wise and greedy search algorithms by fostering a more efficient exploration of solution spaces. Pairwise Preference Reward Model (PPRM), inspired by Reinforcement Learning from Human Feedback (RLHF), is then used to model pairwise preferences between solutions, utilizing an Enhanced Borda Count (EBC) method to synthesize these preferences into a global ranking score to find better answers. This approach addresses the challenges of scoring variability and non-independent distributions in mathematical reasoning tasks. The framework has been tested on general and advanced benchmarks, showing superior performance in terms of search efficiency and problem-solving capability compared to existing methods like ToT and rStar, particularly in complex Olympiad-level benchmarks, including AIME24 and AMC23.
@@ 1 Introduction
Mathematical reasoning represents a great challenge in artificial intelligence, with broad applications across automated theorem proving, mathematical problem solving, and scientific discovery (Ahn et al., 2024). Recently, significant strides have been made by Large Language Models (LLMs) like GPT-4 (Achiam et al., 2023) in general mathematical tasks involving arithmetic or geometric problem-solving (Cobbe et al., 2021; Sun et al., 2024; Ying et al., 2024). However, complex mathematical reasoning remains challenging, especially at the Olympiad-level benchmarks such as AIME (MMA, ). One intuitive way is to break the complex problem-solving solutions into step-by-step thoughts, known as reasoning paths, as shown in Chain-of-Thought (CoT) (Wei et al., 2022). Although prompt-based methods can directly construct such reasoning paths, they remain limited in efficiency and accuracy. As a result, researchers have begun exploring search-based optimization algorithms to enhance the generation of effective reasoning paths.
One line of research focuses on generating solutions based on complete reasoning paths. Tree search techniques with backtracking capabilities are proposed to address the inefficiencies of naive methods, such as Tree-of-Thought (ToT) (Yao et al., 2024) and Breadth-First Search (BFS) based on greedy strategies (Zhang et al., 2024). Heuristic tree search methods like RAP (Hao et al., 2023) and Q* (Wang et al., 2024) treat each step in a reasoning path as a state and the generation of steps as actions, framing the search for reasoning paths as a Markov Decision Process (MDP) problem. Monte Carlo Tree Search (MCTS) methods have further improved the efficiency of path searches in reasoning paths. Yet, the generation of steps in reasoning paths faces challenges due to the infinite action space of thought steps, particularly under limited search conditions, which constrains the performance of step-wise search methods. Unlike methods that generate a single step simultaneously, another approach considers the entire CoT solution as an independent state with iterative optimization actions. This method, known as rewriting, includes techniques such as reflection and self-refine, where optimization of a solution depends on its previous contents and typically involves only local corrections, ensuring that the new solution remains within the “trust region” of the original. This strategy offers the iterative optimization methods a proximal optimization characteristic, making the action space more compact. However, this continuous and greedy iterative approach also suffers from low search efficiency and the risk of local optima, limiting the performance ceiling of rewriting methods.
Another line of research focuses on evaluating the quality of reasoning paths, which includes outcome and process reward models (Uesato et al., 2022). The former focuses on whether the final answer of a reasoning path is correct, while the latter emphasizes the correctness of each step and the logical connections between steps. Although both approaches enable models to score using a scalar, labelled data for training these reward models are scarce. Furthermore, the scoring standards in mathematical reasoning tasks are highly variable, and the reward distribution for different problems is typically not i.i.d. (independently and identically distributed), with each problem presenting unique scoring characteristics. This non-independence complicates the scaling up of single-scalar score-based reward models and their abilities to detail local preference relations between answers.
To address the above challenges, we design the LLaMA-Berry framework (in Figure 1) based on LLaMA-3.1-8B Instruct (Meta, 2024b), which includes two novel methods called Self-Refine applied to Monte Carlo Tree Search (SR-MCTS) and Pairwise Preference Reward Model (PPRM). Specifically, SR-MCTS is a novel Markov Decision Process (MDP) framework focusing on a path-wise problem-solving search where a complete solution is defined as an independent state, and self-refine is considered an optimization action. Then, Monte Carlo Tree Search (MCTS) is leveraged for the exploration of optimal solutions with guiding of PPRM evaluations. PPRM is inspired by Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2017) techniques; we have transformed the task of reward models into an RL problem that models preferences between solutions, then aggregating local preferences into a global ranking score with an Enhanced Borda Count (EBC) method.
Finally, evaluations on general benchmarks such as GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) have demonstrated that our method outperforms baseline approaches like ToT (Yao et al., 2024) and rStar (Qi et al., 2024) in both search efficiency and performance. Moreover, in challenging Olympiad-level benchmarks such as AIME24 and AMC23 and in the Graduate-level GPQA Diamond (Rein et al., 2023) benchmark, our approach has elevated the performance of LLaMA-3.1-8B without any further training to a level comparable to proprietary solutions including GPT-4 Turbo. These encouraging results suggest that our method can effectively improve the LLM’s reasoning ability with only a small amount of data, and this capability can be generalized to other fields, such as physics and chemistry. Furthermore, it can enable multimodal reasoning, allowing the model to achieve even greater reasoning power.
To summarize, our contributions in LLaMA-Berry framework integrate three essential components:
  • SR-MCTS establishes a novel MDP framework that treats complete path-wise solutions as independent states and utilizes Self-Refine as an optimization action, balancing efficiency and performance in search.
  • Inspired by Reinforcement Learning from Human Feedback (RLHF), PPRM leverages the preference relationship between solutions to evaluate their quality, which avoids the volatility of absolute scores while providing a more guided exploration of optimal paths.
  • Finally, EBC method aggregates the local preferences into a global ranking score, allowing the model to identify the overall patterns and enhance the decision-making process.
@@
Figure 1: The main pipeline of LLaMA-Berry, where $S_{i}$ stand for problem-solving solutions and $C_{i}$ stands for critiques. This Pipeline consists of four phases detailed in 2.2, including Selection, Expansion, Evaluation and Backpropagation.
@@ 2.1 Preliminary
In the realm of mathematical problem-solving, we face significant challenges that require systematic approaches to derive accurate and efficient solutions. A core issue in this field is generating and optimizing reasoning paths, which can be formalized as a path-wise Markov Decision Process (MDP) framework. This framework allows us to define the state space $S$, where each state $s$ represents a complete solution to a given problem, and the action space $A$ consists of all feasible rewriting actions $a$ that make transitions between states.
The central issues we aim to address include the development of the value function $Q(s,a)$, which quantifies the expected rewards from executing action $a$ at state $s$, defined as:
@@
Where $T(s,a)$ indicates the transition to another solution $s^{\prime}$ from solution $s$ with rewriting operation denoted as $a$. Our primary objective is to identify the optimal state $s^{*}$, which represents the best solution to the problem. This optimal state can be achieved through the selection of actions that maximize the value function, guiding us toward the most desirable outcomes:
$s^{*}=\arg\max_{s^{\prime}\in S}\mathbb{E}[Q(s)]$ (2)
2.2 Self-refine applied to Monte Carlo Tree Search
Monte Carlo Tree Search (MCTS) is an effective reinforcement learning method within the Markov Decision Processes (MDP) framework, employing states, actions, and cost functions through sampling. The algorithm follows four key steps: selection, expansion, simulation/evaluation, and backpropagation. In the selection phase, the root node is expanded using the Upper Confidence Bound applied to Trees (UCT) algorithm, which selects a node $s$ by balancing exploration and exploitation, that is,
$a=\arg\max_{a\in A(s)}\left(Q(s,a)+c\cdot\sqrt{\frac{\ln N(s)}{N(s,a)}}\right),$ (3)
where $N(s)$ is the visitation count of node $s$, $N(s,a)$ is the action frequency, and $c$ is a parameter controlling exploration. In the expansion phase, node $s$ generates subsequent states $s^{\prime}$, added as new nodes in tree $\mathcal{T}$. The evaluation phase typically uses simulations or heuristics to estimate the Q-values for these nodes. Finally, during backpropagation, the estimated Q-values are updated retroactively from the leaf nodes to the root. This iterative process allows MCTS to explore and exploit the state space effectively, refining decision-making by balancing the exploration of new options with the exploitation of known high-value paths.
@@ 2.2 Self-refine applied to Monte Carlo Tree Search
As shown in Figure 1, the SR-MCTS method combines Monte Carlo Tree Search (MCTS) with Self-Refine, which optimizes the solution search via a structured tree. This method continuously evaluates and improves solutions by combining the iterative nature of MCTS with the self-critique and rewriting capabilities of LLMs as following sections:
@@
Selection Phase: The selection phase identifies a node $s_{i}$ from the search tree $\mathcal{T}$ for expansion, where each node represents a complete solution state. The Upper Confidence Bound applied to Trees (UCT) algorithm is employed to select the optimal node, with dynamic pruning used to avoid local optima. A node $s_{i}$ is considered fully explored when its child nodes reach a predefined limit, and at least one child node’s Q value exceeds the Q value of $s_{i}$.
@@
Expansion Phase: In the expansion phase, as shown in Figure 2, the selected answer $s_{i}$ is expanded by generating successor answers through a Self-Refine process, which includes a Critizing and Rewriting process. The Critizing generates a critique $c_{i}=C(s_{i})$ that identifies drawbacks (e.g., mathematical wrongs or logical fault) in the current chosen answer $S_{i}$, and then Rewriting process generates a new answer $s_{i+1}=R(s_{i},c_{i})$. In practice, to simplify the problem, we assume this process is deterministic, ensuring that the same original state of solutions $s_{i}$ consistently produces the same successor state of solution $s_{i+1}$. The new state of solution $s^{\prime}$ is then added to the search tree $\mathcal{T}$ as a new node.
@@
Evaluation Phase: The evaluation phase calculates the value $Q(s^{\prime})$ of the newly generated node $s^{\prime}$ using the Pairwise Preference Reward Model (PPRM). The evaluation involves two steps: global and local value assessments. The global value $Q_{g}(s^{\prime})$ is determined by the quantile of $s^{\prime}$ in a win-loss matrix $M$, which reflects the win-loss relationships between nodes. The local value $Q_{l}(s^{\prime})$ is derived from comparisons with adjacent nodes in the search tree $\mathcal{T}$. The total value $Q(s^{\prime})$ is then computed as a weighted combination of global and local values: $Q(s^{\prime})=\alpha Q_{g}(s^{\prime})+(1-\alpha)Q_{l}(s^{\prime})$, where $\alpha$ controls the relative influence of each component.
@@
Backpropagation Phase: In the backpropagation phase, the value $Q(s^{\prime})$ of the new node is propagated back to its parent node $s_{i}$, updating Q value of $s_{i}$ as a discounted sum of its child nodes’ Q values: $Q(s_{i})=(1-\gamma)Q(s_{i})+\gamma Q(s^{\prime})$. The discount factor $\gamma$ represents the importance of future rewards. This iterative update mechanism ensures that the values of parent nodes are progressively refined, enhancing the guidance for future selections.
@@
Additionally, to control the growth of the search tree, the SR-MCTS method restricts the maximum number of rollouts $N_{max}$ The search process terminates when the restriction is reached, imposing limits on the unbounded expansion of the tree. The overarching objective of SR-MCTS is to maximize the expected highest $Q$ value of all existing nodes $S$, guiding us towards the most desirable outcomes $s^{*}$, ensuring that the search method efficiently converges to high-quality solutions.
Figure 2: (Left) Self-Refine Process and (Right) Preference Prediction process of PPRM
@@ 2.3 Pairwise Preference Reward Model
In mathematical problem-solving tasks, reliable evaluation of solutions is crucial, as it leads to better estimation of Q-values, thereby offering improved guidance. Existing reward models typically evaluate solutions by giving absolute scores, such as (PRM) (Lightman et al., 2023) and outcome reward mode (ORM) (Yu et al., 2023a). However, the score-based reward models may fall short in leveraging the instruction following capability of LLMs or effectively handling the variations in data distributions and exhibit volatility, especially when the differences between different solutions are subtle. To address this issue, we propose the Pairwise Preference Reward Model (PPRM), which leverages a comprehensive preference dataset incorporating substantial samples from both approaches to learn preference relationships among mathematical solutions. Please refer to Appendix for details about training and inference of PPRM.
For two answers to a given mathematical problem, denoted as $a_{1}\succ a_{2}|\phi$ where $a_{1}$ and $a_{2}$ denote the preferred and dispreferred completion, PPRM predicts their relative quality using a pairwise partial ordering, represented by the following probability formula:
@@
Where $\phi(a_{1}\succ a_{2})$ denotes the feature representation of a partial ordering relation between solution $a_{1}$ and $a_{2}$, with $\phi$ representing the parameters of the model. In our method, $a_{1}\succ a_{2}$ are represented by tokens of the LLM, and $P(a_{1}\succ a_{2}\mid\phi)$ is estimated using the logits value of tokens calculated by the LLM.
Then, inspired by advancements in Language Interface Fine-Tuning (LIFT) (Dinh et al., 2022), we frame the training process of PPRM as a question-answering task. The model is tasked with answering the question, ”For Question $Q$, is solution $a_{1}$ better than solution $a_{2}$?” as shown in Figure 2. To form a robust training objective, the predicted token labels $\hat{y}$ (’Yes’ or ’No’) are evaluated using the indicator function $\mathbf{I}$:
@@
Finally, a pairwise preference dataset $\mathcal{D}$ that contains millions of mathematical problem-solving solution pairs is converted into a dataset $\mathcal{D}^{\prime}$ suitable for a question-answering task. We employ RLHF techniques to train the model to improve its performance in the partial-order prediction question-answering task. Subsequently, the Direct Preference Optimization (DPO) (Rafailov et al., 2024) method is utilized to find the optimal $P_{\phi}$ by maximizing the objective $\arg\max_{\phi}\mathbb{E}_{P}[\mathbf{I}(\hat{y},y)]$.

### v2

# LLaMA-Berry: Pairwise Optimization for Olympiad-level Mathematical Reasoning via O1-like Monte Carlo Tree Search

@@ Abstract
This paper presents an advanced mathematical reasoning framework, LLaMA-Berry, for enhancing the problem-solving ability of large language models (LLMs). The framework combines Monte Carlo Tree Search with Self-Refine (SR-MCTS) to optimize the reasoning paths and utilizes a pairwise reward model to evaluate different paths globally. By leveraging the self-critique and rewriting capabilities of LLMs, our SR-MCTS overcomes the inefficiencies and limitations of conventional step-wise and greedy search algorithms by fostering a more efficient exploration of solution spaces. To guide the search process, we propose Pairwise Preference Reward Model (PPRM) to predict pairwise preferences between solutions through instruction-following capabilities trained by Reinforcement Learning from Human Feedback (RLHF). Finally, the Enhanced Borda Count (EBC) method is adopted to synthesize pairwise preferences into global quantile scores for evaluations. This approach addresses the challenges of scoring variability and non-independent distributions in mathematical reasoning tasks. The framework has been tested on general and advanced benchmarks, showing superior search efficiency and performance compared to existing open-source and closed-source methods, particularly in complex Olympiad-level benchmarks, including AIME24 and AMC23.
LLaMA-Berry: Pairwise Optimization for Olympiad-level Mathematical Reasoning via O1-like Monte Carlo Tree Search
Di Zhang1,2thanks: These authors contributed equally., Jianbo Wu3∗, Jingdi Lei2∗, Tong Che4∗ Jiatong Li5, Tong Xie6, Xiaoshui Huang7, Shufei Zhang2, Marco Pavone8 Yuqiang Li2thanks: Corresponding author , Wanli Ouyang2, Dongzhan Zhou2† 1Fudan University, 2Shanghai Artificial Intelligence Laboratory, 3University of California, Merced 4Independent Researcher, 5Hong Kong Polytechnic University, 6University of New South Wales 7Shanghai Jiao Tong University, 8Stanford University {liyuqiang,zhoudongzhan}@pjlab.org.cn
@@ 1 Introduction
Figure 1: The main pipeline of LLaMA-Berry, where $S_{i}$ stand for problem-solving solutions and $C_{i}$ stands for critiques. The pipeline consists of four phases detailed in Section 2.2, including selection, expansion, evaluation, and backpropagation.
Mathematical reasoning represents a great challenge in artificial intelligence, with broad applications across automated theorem proving, mathematical problem solving, and scientific discovery (Ahn et al., 2024). Recently, significant strides have been made by large language models (LLMs) like GPT-4 (Achiam et al., 2023) in general mathematical tasks involving arithmetic and geometric problem-solving (Cobbe et al., 2021; Sun et al., 2024; Ying et al., 2024). However, complex mathematical reasoning remains challenging, especially at the Olympiad-level benchmarks such as AIME (MAA, 2024).
An intuitive approach to improving problem-solving is to break solutions into step-by-step reasoning paths (Lightman et al., 2023; Luo et al., 2024a), as demonstrated in Chain-of-Thought (CoT Wei et al., 2022). While prompt-based methods can effectively facilitate the construction of such reasoning paths, they may still encounter challenges due to the lack of comprehensive feedback during the generation process, which can affect efficiency (Paul et al., 2023). In contrast to stepwise generation methods, another promising line of research treats the entire solution as an independent state, employing rewriting capabilities to refine the solutions, such as in Self-Refine (Madaan et al., 2023a) and Reflexion (Shinn et al., 2024). However, these approaches, while innovative, may occasionally face challenges like being susceptible to local optima or potentially drifting towards suboptimal solutions due to flawed feedback, which could impact their maximum potential performance.
In addition to generating reasoning paths, effective solution evaluation is crucial, with models like the outcome reward model (ORM) and process reward model (PRM) (Uesato et al., 2022) serving as valuable examples. The ORM focuses on the correctness of the final answer in a reasoning path, while the PRM emphasizes the correctness of each step in the process. While both methods enable reward models to assign scalar scores, obtaining reliable labeled data for training these reward models remains a significant challenge. Moreover, the scoring standards for mathematical reasoning tasks can vary significantly, as each problem presents unique characteristics. This variation complicates the scaling of reward models and hinders their ability to capture local preference relations between solutions. Although trained using language models, these reward models have yet to fully leverage instruction-following capabilities, which may limit their effectiveness in handling more complex reasoning tasks (Zhang et al., 2024a).
To improve the efficiency of solution search in mathematical problems, we treat a complete solution as an independent state and apply Self-Refine to optimize previous solutions in order to obtain better ones. In the Self-Refine process, feedback from critiques is utilized to make the search more efficient compared to stepwise reasoning path generation. Furthermore, we incorporate Monte Carlo Tree Search (MCTS Kocsis and Szepesvári, 2006) to replace the iterative manner in naive Self-Refine, enhancing the solution search. MCTS leverages signals from the evaluation process to assess solutions and uses the Upper Confidence Bound applied to Trees (UCT) method to balance exploration and exploitation. This approach enables the search process to effectively exploit higher-quality solutions and explore those with greater potential for improvement, while avoiding getting trapped in suboptimal local minima.
In evaluation process, utilizing the instruction-following capabilities trained by Reinforcement Learning from Human Feedback (RLHF Christiano et al., 2017), Pairwise Preference Reward Model (PPRM) transforms the absolute rewards calculation into preferences prediction between solutions to calculate rewards. The approach reduces the variability with scoring characteristics and thus leads to a more robust and consistent evaluation of different solutions. To overcome the locality limitations inherent in pairwise comparisons, we employ the Enhanced Borda Count (EBC) method to aggregate local preference evaluations into global quantile scores, leading to more informed decision-making and, ultimately, better solutions. Combining the PPRM and EBC method not only enables the reward model to learn a more robust reward signal but also captures the global characteristics of the solution space, ensuring more reliable comparisons.
Our contributions are summarized as follows: (1) We propose SR-MCTS, a novel Markov Decision Process (MDP) framework that treats entire solutions as states and Self-Refine as optimization action to perform advanced solution search with MCTS. (2) PPRM is developed to leverage the preference relationship between solutions to evaluate their quality, which avoids the volatility of absolute scores while providing a more guided exploration of optimal paths. We adopt the EBC method to convert the local preferences into global evaluations. (3) We verify the effectiveness of LLaMA-Berry on multiple benchmarks, which outperforms baseline approaches like ToT (Yao et al., 2024) and rStar (Qi et al., 2024) in both search efficiency and accuracy. Notably, LLaMA-Berry enhances the performance of LLaMA-3.1-8B, making the 8B-level model comparable to proprietary models, including GPT-4 Turbo on Olympiad-level mathematical reasoning without additional training.
@@ 2.1 Preliminary
One of the core challenges in mathematical problem-solving is to generate and optimize reasoning paths to derive high-quality solutions. We formalize this process in a path-wise Markov Decision Process (MDP) framework, where each state $s$ in the state space $S$ represents a complete solution to a given problem, and the action space $A$ consists of all feasible rewriting actions $a$ that make transitions between states.
Figure 2: Preference prediction process of PPRM and global quantile score based on Enhanced Borda Count method.
In the framework, we aim to quantify the expected reward $Q(s,a)$ from executing action $a$ at state $s$, that is,
@@
, where $T(s,a)$ indicates the transition from $s$ to another solution $s^{\prime}$ via the rewriting action $a$. Our primary objective is to identify the optimal state $s^{*}$ that represents the best solution. We can reach $s^{*}$ by selecting actions that maximize the reward, guiding us toward the most desirable outcome, as demonstrated in Equation 2.
$s^{*}=\arg\max_{s^{\prime}\in S}Q(s^{\prime})$ (2)
2.2 Self-Refine applied to MCTS
@@
As shown in Figure 1, SR-MCTS integrates Monte Carlo Tree Search (MCTS) with the Self-Refine mechanism to continuously evaluate and optimize the solution search. This integration leverages the iterative nature of MCTS and the self-improvement capabilities of LLMs, thereby improving the search outcomes.
Monte Carlo Tree Search (MCTS) is an effective method within the Markov Decision Processes (MDP) framework, employing states, actions, and value functions through sampling. The algorithm follows four key steps: selection, expansion, evaluation, and backpropagation. In the selection phase, the root node is expanded using the Upper Confidence Bound applied to Trees (UCT) algorithm, which selects a node $s$ by balancing exploration and exploitation:
$a=\arg\max_{a\in A(s)}\left(Q(s,a)+c\cdot\sqrt{\frac{\ln N(s)}{N(s,a)}}\right)$ (3)
, where $N(s)$ is the visitation count of node $s$, $N(s,a)$ is the action frequency, and $c$ is a parameter controlling exploration. In the expansion phase, node $s$ generates subsequent states $s^{\prime}$, added as new nodes in the tree $\mathcal{T}$. The evaluation phase typically uses simulations or heuristics to estimate the Q-values for these nodes. Finally, during backpropagation, the estimated Q-values are updated retroactively from the leaf nodes to the root. This iterative process allows MCTS to refine decision-making by balancing the exploration of new paths with the exploitation of known high-value paths.
@@
Selection phase. The selection phase identifies a node $s_{i}$ from the search tree $\mathcal{T}$ for expansion, where each node represents a complete solution state. The Upper Confidence Bound applied to Trees (UCT) algorithm is employed to select the optimal node, with dynamic pruning used to avoid local optima. A node $s_{i}$ is considered fully explored when its child nodes reach a predefined limit, and at least one child node’s Q value exceeds the Q value of $s_{i}$.
@@
Expansion phase. In the expansion phase, as shown in Figure 1, the selected answer $s_{i}$ is expanded by generating successor answers through a Self-Refine process, which includes a Criticizing and Rewriting process. The Criticizing process generates a critique $c_{i}=C(s_{i})$ that identifies drawbacks (e.g., mathematical wrongs or logical faults) in the current chosen answer $S_{i}$, and then Rewriting process generates a new answer $s_{i+1}=R(s_{i},c_{i})$. In practice, to simplify the problem, we assume this process is deterministic, ensuring that the same original state of solutions $s_{i}$ consistently produces the same successor state of solution $s_{i+1}$. The new state of solution $s^{\prime}$ is then added to the search tree $\mathcal{T}$ as a new node.
@@
Evaluation phase. The evaluation phase calculates the value $Q(s^{\prime})$ of the newly generated node $s^{\prime}$ using the Pairwise Preference Reward Model (PPRM). The evaluation involves two steps: global and local value assessments. The global value $Q_{g}(s^{\prime})$ is determined by the quantile of $s^{\prime}$ in a win-loss preference matrix $\mathbf{M}$, which reflects the win-loss relationships between nodes. The local value $Q_{l}(s^{\prime})$ is derived from comparisons with adjacent nodes in the search tree $\mathcal{T}$. The total value $Q(s^{\prime})$ is then computed as a weighted combination of global and local values: $Q(s^{\prime})=\alpha Q_{g}(s^{\prime})+(1-\alpha)Q_{l}(s^{\prime})$, where $\alpha$ controls the relative influence of each component.
@@
Backpropagation phase. In the backpropagation phase, the value $Q(s^{\prime})$ of the new node is propagated back to its parent node $s_{i}$, updating Q value of $s_{i}$ as a discounted sum of its child nodes’ Q values: $Q(s_{i})=(1-\gamma)Q(s_{i})+\gamma Q(s^{\prime})$. The discount factor $\gamma$ represents the importance of future rewards. This iterative update mechanism ensures that the values of parent nodes are progressively refined, enhancing the guidance for future selections.
@@
Additionally, to control the growth of the search tree, the SR-MCTS method restricts the maximum number of rollouts $N_{max}$ The search process terminates when the restriction is reached, imposing limits on the unbounded expansion of the tree. The overarching objective of SR-MCTS is to maximize the expected highest $Q$ value of all existing nodes $S$, guiding us towards the most desirable outcome $s^{*}$, ensuring that the search process efficiently converges to high-quality solutions.
@@ 2.3 Pairwise Preference Reward Model
Reliable evaluation of different solutions plays a crucial role in mathematical problem-solving tasks as it leads to better estimation of Q-values, thereby offering better guidance. Existing reward models typically evaluate solutions by giving absolute scores, such as process reward model (PRM Lightman et al., 2023) and outcome reward model (ORM Yu et al., 2023). However, the score-based reward models may fall short in leveraging the instruction-following capability of LLMs or effectively handling the variations in scoring standards, especially when the differences between solutions are subtle. To address this issue, we propose the Pairwise Preference Reward Model (PPRM), which leverages a comprehensive preference dataset incorporating substantial samples from both PRM and ORM approaches (Toshniwal et al., 2024; Lightman et al., 2023) to learn preference relationships among mathematical solutions.
For two solutions ($a_{1}$ and $a_{2}$) to a given mathematical problem, we use $a_{1}\succ a_{2}$ to represent the situation where $a_{1}$ is preferred over $a_{2}$. PPRM predicts their relative quality using a pairwise partial ordering, represented by the following probability formula:
@@
, where $P(a_{1}\succ a_{2}\mid\phi)$ denotes the probability of a partial ordering relation between solution $a_{1}$ and $a_{2}$, with $\phi$ representing the parameters of the reward model. In our method, $a_{1}\succ a_{2}$ are represented by tokens of an LLM, and $P(a_{1}\succ a_{2}\mid\phi)$ is estimated using the logits value of tokens calculated by the LLM.
Then, inspired by advancements in Language Interface Fine-Tuning (LIFT Dinh et al., 2022), we frame the training process of PPRM as a question-answering task to leverage the instruction-following capabilities of LLMs. The model is tasked with answering the question, "For Question $Q$, is solution $a_{1}$ better than solution $a_{2}$?" as shown in Figure 2. To form a robust training objective, the predicted token labels $\hat{y}$ (’Yes’ or ’No’) are evaluated with ground truth label $y$ using the indicator function $\mathbf{I}$:
@@
Finally, a pairwise preference dataset $\mathcal{D}$ that contains millions of mathematical problem-solving solution pairs is converted into a dataset $\mathcal{D}^{\prime}$ suitable for a question-answering task. We employ RLHF techniques to train the model to improve its performance in the partial-order prediction question-answering task. Subsequently, the Direct Preference Optimization (DPO Rafailov et al., 2024) method is utilized to find the optimal $P_{\phi}$ by maximizing the objective $\arg\max_{\phi}\mathbb{E}_{P}[\mathbf{I}(\hat{y},y)]$. Please refer to Appendix C for details about training and inference of PPRM.